OpenStudy (anonymous):

assuming U, V and W are vectors, prove that (UxV)*[(VxW)x(WxU)]=[U*(VxW)]^2, where x is cross product, * is dot product.

OpenStudy (anonymous):

i know Ax(BxC)=B(A.C)-C(A.B), and A.(BxC)=B.(CxA)=C.(AxB), but i dont know where to start from.

OpenStudy (amistre64):

im not seeing the v u ws right, they tend to merge into one .....

OpenStudy (anonymous):

ugh o.o?

OpenStudy (amistre64):

i assume we have to do it with identities then?

OpenStudy (anonymous):

the question hinted to use Ax(BxC)=B(A*C)-C(A*B), and A*(BxC)=B*(CxA)=C*(AxB), but i cant see where's the catch

OpenStudy (amistre64):

is the dot product distributive?

OpenStudy (amistre64):

a.(bxc) = a.b ? a.c

OpenStudy (amistre64):

dot produce a scalar, and crossing a scalar is undefined so i really got to read up on these properties again :)

OpenStudy (anonymous):

er yes dot is distributive and commutative, scalar is not commutative and associative

OpenStudy (anonymous):

scalar is not commutative and not* associative

OpenStudy (anonymous):

erm sry not scalar, its cross, lol

OpenStudy (amistre64):

\[\vec a.(\vec b\times \vec c)=\vec a\vec c-\veca \vec b \] oy this thing is a headache :)

OpenStudy (anonymous):

er izit possible to let UxV be A, VxW be B, and WxU be C,

OpenStudy (amistre64):

(UxV) * ( (VxW)x(WxU) ) = ( U*(VxW) )^2 (WxU)*((UxV)x (VxW)) - (VxW) ( (UxV)x(WxU) ) = ( U*(VxW) )^2 yes, its prolly best to redifine them so that he mess cleans up; since all they produce are another vector

OpenStudy (amistre64):

(UxV) * ( (VxW)x(WxU) ) = ( U*(VxW) )^2 A.(BxC) = (U.B)^2 the thing is now we have to define "U" from the others in order to pair it up in the end

OpenStudy (amistre64):

but once you get thist o something close in resemblance; you can switch back to the UVW mess and see if it pans out

OpenStudy (anonymous):

erm does (VxW)x(WxU) = VxWxWxU = VxWxU?

OpenStudy (anonymous):

kinda mixed up with boolean already LOL

OpenStudy (amistre64):

A.(BxC) = C.(AxB) - B.(AxC) WxW is only 1, if W = W^-1

OpenStudy (amistre64):

in general, WxW not= W perse

OpenStudy (amistre64):

and since there is no indication that these are perp to each other we should keep it as general as possible

OpenStudy (amistre64):

just my thoughts about it tho

OpenStudy (anonymous):

ouh i'll try again.

OpenStudy (amistre64):

i cant parse it with any degree of certainty :) so i wish you the best of luck on it; and when all else fails, google it :)

OpenStudy (anonymous):

cant google it, lol. maximum only managed to obtain the 2 properties as given in the question。

OpenStudy (anonymous):

anyway thx for your opinion ;D